OpenStudy (anonymous):

Find the sale price if the price of $140 is discounted 25%.

OpenStudy (saifoo.khan):

25 x 140/100

OpenStudy (saifoo.khan):

$35

OpenStudy (anonymous):

25% of $140=140*25/100=140/4=$35 selling price= actual price- discount=140-35=105

OpenStudy (saifoo.khan):

therefore: 140 - 35 = 105

OpenStudy (anonymous):

Thank you!

OpenStudy (saifoo.khan):

uw